OpenStudy (anonymous):

Write y = -4x^2 - 16x - 14 in vertex form A. y = -4(x+2)^2+2 B. y = 4(x+4)^2+2 C. y = -4(x-16)^2-14 D. y = (x+2)^2 - 14 I think it's either A or B, but I don't know how to work these...?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y = -4x^2 - 16x - 14 }\) your first step would be: \(\large\color{black}{ \displaystyle y = -4(x^2 + 4x) - 14 }\)

OpenStudy (solomonzelman):

do you see what I did in the first step?

OpenStudy (anonymous):

(h,k) is vertex to find h http://mathbitsnotebook.com/Algebra1/Quadratics/vertex1.gif

OpenStudy (solomonzelman):

we need to write the equation though, Austin...

OpenStudy (anonymous):

I'm not sure what you did to get the 4x ?

OpenStudy (solomonzelman):

I factored the first 2 terms out of -4.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y = -4x^2-16x - 14 }\) \(\large\color{black}{ \displaystyle y = \color{blue}{-4}(x^2) \color{blue}{-4}( 4x) - 14 }\) \(\large\color{black}{ \displaystyle y = -4(\color{green}{x^2 + 4x}) - 14 }\)

OpenStudy (solomonzelman):

did this make sense just now?

OpenStudy (anonymous):

Okay, yeah... Then would you add 14 to both sides?

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

we got till here so far: \(\large\color{black}{ \displaystyle y = -4(\color{green}{x^2 + 4x}) - 14 }\) what number would you add to \(\large\color{black}{ \displaystyle \color{green}{x^2 + 4x}}\) to make it a perfect square trinomial? \(\large\color{black}{ \displaystyle \color{green}{x^2 + 4x}+?}\)

OpenStudy (anonymous):

Oh... I have absolutely no idea...